r is 9 percent, RR are $10 million, and (total) reserves (R) are $50 million.
What do excess reserves (ER) equal? What do DD equal?
1
Expert's answer
2020-04-10T09:12:40-0400

You can calculate a bank's excess reserves, if any, by using the following formula:

"EX = LR-RR"

Where:

EX - excess reserves 

LR - legal reserves 

RR - required reserves

"ER=50-10=40"

"10=0.09 * DD"

Where 0.09=9%

"DD=10\/0.09=111.11"

DD = 111.11 million $


https://study.com/academy/lesson/excess-reserves-definition-formula-quiz.html